# Continual Memorization of Factoids in Large Language Models

### v1

@@ Abstract
Large language models (LLMs) can absorb a massive amount of knowledge through pretraining, but pretraining is inefficient for acquiring long-tailed or specialized facts. Therefore, fine-tuning on specialized or new knowledge that reflects changes in the world has become popular, though it risks disrupting the model’s original capabilities. We study this fragility in the context of continual memorization, where the model is trained on a small set of long-tail factoids (subject-relation-object associations) and must retain these factoids after multiple stages of subsequent training on other datasets. Continual memorization focuses on the specific challenge of retaining long-tail factoids, whereas general continual learning aims to maintain the LLM’s capabilities across a wide range of generic tasks (e.g., reasoning, commonsense knowledge). Through extensive experiments, we show that LLMs suffer from forgetting across a wide range of subsequent tasks, and simple replay techniques do not fully prevent forgetting, especially when the factoid datasets are trained in the later stages. We posit that there are two ways to alleviate forgetting: 1) protect the memorization process as the model learns the factoids, or 2) reduce interference from training in later stages. With this insight, we develop an effective mitigation strategy: REMIX (Random and Generic Data Mixing). REMIX prevents forgetting by mixing generic data sampled from pretraining corpora or even randomly generated word sequences during each stage, despite being unrelated to the memorized factoids in the first stage. REMIX can recover performance from severe forgetting, often outperforming replay-based methods that have access to the factoids from the first stage. We then analyze how REMIX alters the learning process and find that successful forgetting prevention is associated with a pattern: the model stores factoids in earlier layers than usual and diversifies the set of layers that store these factoids. The efficacy of REMIX invites further investigation into the underlying dynamics of memorization and forgetting, opening exciting possibilities for future research.
11footnotetext: Equal contribution. Code available at: URL.
@@ 1 Introduction
Figure 1: The continual memorization setting. In stage 1 (red box), a pretrained model is trained to convergence on a factoid dataset $D_{A}$ to yield the fine-tuned model $\theta_{A}$. In stage 2, model $\theta_{A}$ is further trained on either a factoid dataset or a non-factoid dataset (blue box) to obtain model $\theta_{B}$. The final model $\theta_{B}$ is evaluated on the training examples $D_{A}$ in stage 1. REMIX mixes either random word sequences or pretraining data (denoted by $D_{M}$) into training during stages 1 and 2 to prevent forgetting. We describe a two-stage setting in this figure and use it as the basis for most of our experiments, but the general setting naturally extends to multiple stages.
Large language models (LLMs) have shown a remarkable ability to absorb a massive amount of knowledge through large-scale pretraining (Petroni et al., 2019; AlKhamissi et al., 2022; Cohen et al., 2023). Despite their familiarity with common knowledge, they still struggle to capture the long tail (Kandpal et al., 2023). Recent work explains that during the pretraining phase, each piece of knowledge requires many exposures and diverse manifestations to be properly acquired (Allen-Zhu & Li, 2024a; b; Chang et al., 2024).
A straightforward alternative is to finetune the model on a small, domain-specific dataset. However, finetuning on long-tail knowledge can cause unintentional harm by decreasing factuality and exacerbating hallucination (Kang et al., 2024; Gekhman et al., 2024; Zhang et al., 2024; Ghosal et al., 2024). In this regard, the finetuning process bears some similarity to continual learning (McCloskey & Cohen, 1989; Ratcliff, 1990), where one tries to successively train a model on a series of tasks without forgetting earlier ones. Prior research on continual learning in LLMs focuses on general capabilities such as reasoning (Luo et al., 2023a), or broad proxies like the language modeling loss over a general corpus (Yıldız et al., 2024).
In this work, we focus on the challenges unique to the continual learning of factoids – atomic facts representable as subject-object relations. We formalize this setting as continual memorization (Figure 1), where a model is first trained on a small collection of factoids (stage 1), and then must retain their knowledge after training on other datasets (stage 2). We first study how different tasks affect forgetting when placed in the second stage. We find that the effect is maximum for factoid datasets (i.e., datasets consisting of factoids) and that it is less pronounced for non-factoid datasets such as those involving coding, math, or chat abilities. Even more worryingly, we find that typical replay methods, which typically work well for general continual learning, fail to prevent model forgetting when the second stage involves a factoid dataset.
How can an LLM prevent the forgetting of factoids? We intuit that this question may be approached in two ways: 1) teach the model to protect learned knowledge better in the first stage, or 2) reducing the interference of the second stage by manipulating the data distribution. Based on this hypothesis, we propose REMIX (Random and Generic Data Mixing), which combines both approaches. First, REMIX mixes random or generic data into the factoids in the first stage. While surprising at first glance, including a broad range of mixed data teaches the model to diversify where it stores the knowledge – as we show in later analysis of REMIX. This diversification allows it to better protect learned knowledge. In the second stage, jointly learning the mixing data and the stage 2 data avoids overfitting to a narrow distribution, alleviating the negative interference on the learned factoids.
Our experiments demonstrate that REMIX is highly effective at helping the model retain learned factoids: in the most severe case, REMIX increases post-phase 2 accuracy from $13.5\%$ to $53.2\%$. In comparison, replay can only reach $41.6\%$ despite using $10\%$ of the factoids from stage 1. These benefits are seen consistently across several choices of factoid and non-factoid tasks in stage 2. We finally perform a careful analysis of REMIX through Logit Lens (nostalgebraist, 2020) and ablation studies. We find that REMIX teaches the model to both store facts in relatively earlier layers (as opposed to the unmixed case) and diversify their storage to many layers.
@@
  • We formalize the setting of continual memorization, identify its unique challenge of memorizing factoid data, and demonstrate that it cannot be easily addressed with replay.
  • We propose REMIX, a simple strategy that does not require access to the factoids from prior stages; we establish through experiments that REMIX helps models remember factoids better – often increasing accuracy by as much as 3×3$\times$3 ×.
  • Through careful analysis and ablation studies, we find that REMIX operates by teaching the model to protect factoids via diversification and by reducing the negative interference from the later training stages.
@@ 2.1 Problem Definition
We define a factoid to be a triple (subject, relation, object). A dataset $D\in\mathcal{D}$ in this paper is a set of (prompt, response) pairs. A factoid dataset $D\in\mathcal{D}_{\text{fact}}\subset\mathcal{D}$ is a set of factoids formatted as pairs $(\textit{prompt}\equiv(\textit{subject},\textit{relation}),\textit{response}\equiv\textit{object})$. Here, the $\equiv$ sign signifies that the relation instances are formatted using a template defined for each task (e.g., “The <X> of <Y> is” $\rightarrow$ <Z>). If $D\in\mathcal{\mathcal{missing}}D\setminus\mathcal{D}_{\text{fact}}$, we call $D$ a non-factoid dataset.
A language model is a set of parameters $\theta$ that define a mapping $f(\cdot;\theta):x\mapsto P(y)$ from a prompt $x$ to a distribution over responses $y$. We will overload $f(x;\theta)$ to also refer to the response with maximum probability. Given a model $\theta$ and dataset $D$, we denote by $\mathcal{L}(\theta;D)\in\mathbb{R}^{+}$ the loss and $\mathcal{A}(\theta;D)\in[0,1]$ the average exact-match accuracy. We define a factoid $x$ to be familiar to $\theta$ if $\mathcal{A}(\theta;\{x\})=1$ and unfamiliar otherwise. An unfamiliar dataset consists entirely of unfamiliar facts.
@@
We now describe the setting of continual memorization, which consists of two or more stages. We describe the setting with two stages below. Let $D_{A}\in\mathcal{D}_{\text{fact}}$ be a factoid dataset, and $D_{B}\in\mathcal{D}$ be another dataset (factoid or non-factoid). In the first stage, a pretrained model $\theta_{0}$ is trained on $D_{A}$ until convergence to obtain the trained model $\theta_{A}$ with near-zero loss $\mathcal{L}(\theta_{A};D_{A})\approx 0$ and accuracy $\mathcal{A}(\theta_{A};D_{A})\approx 1$. In the second stage, $\theta_{A}$ is further trained on $D_{B}$ until convergence. The resulting model $\theta_{B}$ is evaluated on $D_{A}$ to gauge its retention $\mathcal{A}(\theta_{B},D_{A})$. In this paper, we consider the case where all factoid datasets (in the first as well as second stage—if applicable) are unfamiliar and we refer to them simply as factoid datasets. Typically, one observes $\mathcal{A}(\theta_{B},D_{A})\ll\mathcal{A}(\theta_{A},D_{A})$ due to catastrophic forgetting. Figure 1 illustrates this setting.
@@ 2.2 Constructing Factoid Datasets
We consider a variety of (unfamiliar) factoid datasets in our experiments. These datasets are either 1) constructed synthetically to ensure that they were not seen by the model $\theta_{0}$ during pretraining—such as by generating random key-value mappings, or 2) by filtering factoid datasets to remove familiar instances (more details in § B.4). We further describe the specific choice of datasets for the two stages below.
@@
  • Key-Value Recall (KVR). We construct the Key-Value Recall task by randomly generating $2,000$ unique key-value pairs. Each key and value string contains $8$ characters from the mix of alphabets and number digits. See Figure 1 for an example.
  • PopQA. PopQA (Mallen et al., 2023) is a collection of $14$k questions about long-tail entities. It contains $16$ diverse relationship types presented in a knowledge triplet format sourced from Wikidata (Vrandečić & Krötzsch, 2014). We randomly select $2,000$ unfamiliar data points from the dataset for finetuning.
  • TriviaQA. TriviaQA (Joshi et al., 2017) is a question answering dataset containing over $650$k question-answer pairs. We randomly select $2,000$ unfamiliar examples to fine-tune on.
@@
  • Factoid datasets: LAMA (Petroni et al., 2019), Entity Questions (Sciavolino et al., 2021), WebQA (Berant et al., 2013). In addition, we also explore adding new (and unfamiliar) examples from the distribution of $D_{A}$ (i.e., the same task as in stage 1)–referred to as the “In-Domain” (ID) datasets in our results.
  • Non-factoid datasets: UltraChat (Ding et al., 2023), EvolCode (Luo et al., 2023b), APPS (Hendrycks et al., ), GSM8K (Cobbe et al., 2021), and MATH (Hendrycks et al., 2021), . These datasets exemplify some common non-factoid datasets used for fine-tuning: chat, code and math.
@@
We use Llama-3-8B (Dubey et al., 2024) and Mistral-7B (Jiang et al., 2023) to initialize $\theta_{0}$ in our experiments (both are base models). All of our experiments use the Tulu-v2 prompt template (Ivison et al., 2023), i.e., "<user>...<assistant>..." for both stages. We provide training details in §B.4. Our accuracies are computed as Exact String Match and normalized to $[0,100]$ for all the experiments, as the tasks generally only need to generate a few tokens. We report averaged accuracy across $3$ runs.
@@ 3.1 Understanding the Forgetting Patterns
Factoid Non-Factoid
Stage 1 ID LAMA EQ WQ Avg GSM8K MATH EC APPS UC Avg
KVR   0.5   2.1 17.4 33.8 13.5 24.4 27.3 49.5 26.7 66.6 38.9
PopQA 49.8   7.7 57.8 72.5 47.0 19.0 92.4 77.0 55.1 48.5 58.4
TriviaQA 45.6   4.3 40.5 68.6 39.8   9.4 87.6 54.4 70.4 67.6 57.9
@@
We first establish the forgetting patterns in continual memorization by examining which intervening tasks affect the final accuracy most severely when trained on in the second stage. Table 1 shows the performance degradation of stage 1 tasks after training on stage 2 tasks. We observe that forgetting is most severe when stage 2 is also a factoid dataset, degrading accuracy for Key-Value Recall to $13.5\%$, PopQA to $47.0\%$, and TriviaQA to $39.8\%$ on average. In fact, with LAMA these accuracies fall to $2.1\%,7.7\%$ and $4.3\%$ respectively—far below the numbers seen with non-factoid datasets. This corroborates findings from the continual learning literature which suggest catastrophic forgetting happens when two tasks are similar and therefore interfere (Farajtabar et al., 2020; Bennani et al., 2020; Doan et al., 2021). In general, non-factoid datasets see a lesser effect, though some datasets like GSM8k still bring about a significant drop.
Figure 2: Replay results averaged across all $D_{B}$ for four mixing ratios.

### v2

# Continual Memorization of Factoids in Language Models

@@ Abstract
As new knowledge rapidly accumulates, language models (LMs) with pretrained knowledge quickly become obsolete. A common approach to updating LMs is fine-tuning them directly on new knowledge. However, recent studies have shown that fine-tuning for memorization may be ineffective in storing knowledge or may exacerbate hallucinations. In this work, we introduce a setting we call continual memorization, where a model must memorize and retain a set of factoids through multiple stages of fine-tuning on subsequent datasets. We characterized the forgetting patterns through extensive experiments and show that LMs widely suffer from forgetting, especially when needing to memorize factoids in the second stage. We posit that forgetting can be alleviated by modifying training dynamics: (1) protecting the memorization process when learning factoids or (2) reducing interference from subsequent training stages. Intriguingly, we find that mixing randomly generated word sequences or generic data sampled from pretraining corpora at different training stages effectively mitigates forgetting (REMIX: Random and Generic Data Mixing). REMIX can recover performance from severe forgetting, outperforming replay methods and other continual learning baselines. We analyze how REMIX influences the learning process and find that robust memorization follows a distinct pattern: the model stores factoids in earlier layers than usual and diversifies the layers that retain them, which results in easier recall and manipulate of the learned factoids.111 Code and data available at
https://github.com/princeton-nlp/continual-factoid-memorization.
Machine Learning, ICML
Figure 1: The continual memorization setting. In stage 1 (red box), a pretrained model $\mathcal{M}_{0}$ is trained to convergence on a factoid dataset $D_{A}$ to obtain model $\mathcal{M}_{A}$. In stage 2, model $\mathcal{M}_{A}$ is further trained on either a factoid dataset or a non-factoid dataset (blue box) to obtain model $\mathcal{M}_{B}$. The final model $\mathcal{M}_{B}$ is evaluated on the training examples $D_{A}$ in stage 1. REMIX: mixing random words and pretraining data into training during stages 1 and 2 alleviates forgetting.
@@ 1 Introduction
Language models (LMs) have shown a remarkable ability to absorb massive amounts of knowledge through large-scale pretraining (Petroni et al., 2019; AlKhamissi et al., 2022; Cohen et al., 2023). However, knowledge is not static—new information accumulates quickly while old knowledge becomes obsolete. This dynamic nature necessitates frequent model updates, making costly pretraining impractical. A common approach is to fine-tune the model directly on new knowledge. However, recent studies have shown that fine-tuning is brittle: training on long-tail knowledge can lead to unintended disruptions, such as decreased factuality and exacerbated hallucinations (Kang et al., 2024; Gekhman et al., 2024; Zhang et al., 2024). Furthermore, a fine-tuned model might not properly memorize knowledge, failing to recall or manipulate it effectively (Allen-Zhu & Li, 2024a, b).
In this work, we investigate the dynamics of memorization in a continual learning setting (McCloskey & Cohen, 1989; Ratcliff, 1990), in which the model acquires new information incrementally, one set at a time. While prior research on continual learning in LMs has focused on general capabilities such as reasoning (Luo et al., 2023a) or broad proxies like language modeling loss over a general corpus (Yildiz et al., 2024), we use this framework to study the memorization dynamics of LMs through fine-tuning. We formalize this as continual memorization, in which a model is first trained on a small collection of factoids (factual associations) and must retain this knowledge after training on additional datasets in a subsequent stage. Specifically, we train models to memorize factoid datasets (stage 1) and then evaluate how well these factoids are retained after a second stage of training on a different dataset (Figure 1). We study 2 stages in our main setting, and provide analysis of more stages in §5.
We conduct extensive experiments to characterize forgetting patterns in continual memorization. First, we find that the most severe forgetting occurs when the second stage of training involves another factoid dataset, regardless of whether the facts overlap with those from stage 1. For example, accuracy on TriviaQA drops from $100\%$ (after stage 1) to $39.8\%$ after further training on other factoid datasets, such as LAMA. Forgetting is less pronounced when fine-tuning on non-factoid datasets, such as those involving coding, math, or chat. We also find that long-tail data is the hardest to retain—a randomly generated key-value string is most susceptible to forgetting, with accuracy dropping to $13\%$ after further training on another factoid dataset. This aligns with recent findings that memorizing knowledge (e.g., factoids) causes greater disruption to other model capabilities, which, in our setting, results in more severe forgetting (Kang et al., 2024). Finally, we observe that common experience replay methods, which mix a fraction of data from earlier training stages, fail to prevent forgetting when the second stage involves a factoid dataset, in contrast to their effectiveness in general continual learning settings. For instance, even when mixing in $10\%$ of the factoids from stage 1, the model fails to recover performance beyond $60\%$.
Next, we investigate whether data mixing strategies can mitigate forgetting. Through theoretical derivations, we develop intuition that this question may be approached in two ways: 1) teaching the model to protect learned knowledge better in the first stage, or 2) reducing the interference of the second stage by manipulating the data distribution. Based on this hypothesis, we examine a range of data mixing strategies at each stage. Intriguingly, we find that mixing in either generic pretraining data or even random word sequences leads to a considerable reduction in forgetting. We combine both strategies, and refer to this mitigation as REMIX (Random and Generic Data Mixing). Our experiments demonstrate that REMIX is highly effective at helping the model retain learned factoids: in the most severe case, REMIX increases post-phase 2 accuracy from $13.5\%$ to $53.2\%$. In comparison, replay can only reach $41.6\%$ despite using $10\%$ of the factoids from stage 1. Other common continual learning methods also fall short, where weight regularization (EWC) (Kirkpatrick et al., 2017) and behavior regularization (Sun et al., 2020) both lag behind REMIX. These benefits are seen consistently across several choices of factoid and non-factoid tasks in stage 2.
To understand why these mixing strategies help reduce forgetting, we analyze REMIX using Logit Lens (nostalgebraist, 2020) and ablation studies. Our analysis suggests that including a broad range of mixed data encourages the model to store facts in relatively earlier layers (compared to in the baseline setting), as well as to diversify where it stores the knowledge. This diversification allows it to better protect learned knowledge in subsequent stages of training. In the second stage, jointly learning the mixing data and the stage 2 data helps prevent overfitting to a narrow distribution, alleviating the negative interference on the learned factoids. Finally, we show that more robustly memorized factoids are not only better retained and recalled, but are more easily extracted for manipulation.
@@
  • We formalize the continual memorization setting to study the fragility and remedy of the factoid memorization process in LMs, and demonstrate that it cannot be easily addressed with replay.
  • We find that mixing random and generic data (REMIX) in different stages can greatly mitigate forgetting without accessing the factoids from prior stages.
  • We find that REMIX diversifies the layers where the learned knowledge is stored and tend to store it in earlier layers than the models that suffer from forgetting, shedding light on the patterns of robust memorization.
@@ 2.1 Problem Definition
We define a factoid to be a triplet (subject, relation, object). A dataset $D\in\mathcal{D}$ in this paper is a set of (prompt, response) pairs. A factoid dataset $D\in\mathcal{D}_{\text{fact}}\subset\mathcal{D}$ is a set of factoids formatted as pairs (e.g., “prompt = The <relation> of <subject> is” $\rightarrow$ response = <object>). If $D\in\mathcal{\mathcal{missing}}D\setminus\mathcal{D}_{\text{fact}}$, we call $D$ a non-factoid dataset.
A language model $\mathcal{M}$: $p_{\theta}(y\mid x)$ parameterized by $\theta$ defines a distribution over response $y$ given the prompt $x$. Given a model $\mathcal{M}$ and dataset $D$, we denote by $\mathcal{L}(\theta;D)\in\mathbb{R}^{+}$ the loss and $\mathcal{A}(\mathcal{M};D)\in[0,1]$ the average exact-match accuracy. We define a factoid $x$ to be familiar to $\mathcal{M}$ if $\mathcal{A}(\mathcal{M};\{x\})=1$ and unfamiliar otherwise. An unfamiliar dataset consists entirely of unfamiliar facts.
@@
We now describe the setting of continual memorization, which consists of two or more stages. We describe the setting with two stages below. Let $D_{A}\in\mathcal{D}_{\text{fact}}$ be a factoid dataset, and $D_{B}\in\mathcal{D}$ be another dataset (factoid or non-factoid). In the first stage, a pretrained model $\mathcal{M}_{0}$ is trained on $D_{A}$ until convergence to obtain the trained model $\mathcal{M}_{A}$ with near-zero loss $\mathcal{L}(\theta_{A};D_{A})\approx 0$ and accuracy $\mathcal{A}(\mathcal{M}_{A};D_{A})\approx 1$. In the second stage, $\mathcal{M}_{A}$ is further trained on $D_{B}$ until convergence. The resulting model $\mathcal{M}_{B}$ is evaluated on $D_{A}$ to gauge its retention $\mathcal{A}(\mathcal{M}_{B},D_{A})$. In this paper, we consider the case where all factoid datasets (in the first as well as second stage—if applicable) are unfamiliar and we refer to them simply as factoid datasets. Typically, one observes $\mathcal{A}(\mathcal{M}_{B},D_{A})\ll\mathcal{A}(\mathcal{M}_{A},D_{A})$ due to catastrophic forgetting. Figure 1 illustrates this setting.
@@ 2.2 Constructing Factoid Datasets
We consider a variety of (unfamiliar) factoid datasets in our experiments. These datasets are either 1) constructed synthetically to ensure that they were not seen by the model $\mathcal{M}_{0}$ during pretraining—such as by generating random key-value mappings, or 2) by filtering factoid datasets to remove familiar instances (details in § B.7). We further describe the specific choice of datasets for the two stages below.
@@
Key-Value Recall (KVR): we generate $2,000$ unique key-value pairs, each contains $8$ characters from the mix of alphabets and number digits. PopQA: we sample $2,000$ unfamiliar knowledge triplets from a set of diverse questions and relationships about long-tail entities (Mallen et al., 2023). TriviaQA: we sample $2,000$ unfamiliar question-answer pairs from the dataset  (Joshi et al., 2017). See examples in Figure 1 and §B.8.
@@
  • Factoid datasets: LAMA (Petroni et al., 2019), Entity Questions (Sciavolino et al., 2021), WebQA (Berant et al., 2013). In addition, we also explore adding new (and unfamiliar) examples from the distribution of $D_{A}$ (i.e., the same task as in stage 1) referred to as the “In-Domain” (ID) datasets in our results.
  • Non-factoid datasets: UltraChat (Ding et al., 2023), EvolCode (Luo et al., 2023b), APPS (Hendrycks et al., ), GSM8K (Cobbe et al., 2021), and MATH (Hendrycks et al., 2021b). These datasets exemplify common non-factoid datasets used for finetuning: chat, code and math.
@@
We use Llama-3-8B (Dubey et al., 2024) and Mistral-7B (Jiang et al., 2023) to initialize $\mathcal{M}_{0}$ in our experiments (both are base models). All of our experiments use the Tulu-v2 prompt template (Ivison et al., 2023), i.e., "<user>...<assistant>..." for both stages. We provide training details in §B.7. Our accuracies are computed as Exact String Match and normalized to $[0,100]$ for all the experiments, as the tasks generally only need to generate a few tokens. We report averaged accuracy across $3$ runs.
@@ 3.1 Understanding the Forgetting Patterns
$D_{B}$: Factoid
$D_{A}$ ID LAMA EQ WQ Avg
KVR 0.5 2.1 17.4 33.8 13.5
PopQA 49.8 7.7 57.8 72.5 47.0
TriviaQA 45.6 4.3 40.5 68.6 39.8
$D_{B}$: Non-Factoid
$D_{A}$ GSM8K MATH EC APPS UC Avg
KVR 24.4 27.3 49.5 26.7 66.6 38.9
PopQA 19.0 92.4 77.0 55.1 48.5 58.4
TriviaQA 9.4 87.6 54.4 70.4 67.6 57.9
@@
Figure 2: Intuition behind each mixing strategy. In general, forgetting occurs when $\nabla\mathcal{L}(\theta;D_{A})^{T}\nabla\mathcal{L}(\theta;D_{B})<0$ (angle between red and blue arrows larger than 90 degree). The model goes from $\theta_{0}$ to $\theta_{A}$ in stage 1 (gray arrow), and arrives at $\theta_{B}$ in stage 2 (blue arrow). The translucent blobs represent low-loss region for each dataset. No Mixing: the opposing angle between the red and blue arrows contributes to forgetting. Mixing at Stage 1: the mixing data $D_{M}$ protects memorization by shifting the model parameters to reduce the angle between the red and blue arrows while converging to a low loss on $D_{A}$. Mixing at Stage 2: mixing data $D_{M}$ reduces the interference of $D_{B}$ by lowering the angle between blue and red arrows.
We first establish the forgetting patterns in continual memorization by examining which intervening tasks affect the final accuracy most severely when trained on in the second stage. Table 1 shows the performance degradation of stage 1 tasks after training on stage 2 tasks. We observe that forgetting is most severe when stage 2 is also a factoid dataset, degrading accuracy for KVR to $13.5\%$, PopQA to $47.0\%$, and TriviaQA to $39.8\%$ on average. In fact, with LAMA these accuracies fall to $2.1\%,7.7\%$ and $4.3\%$ respectively—far below the numbers seen with non-factoid datasets. This corroborates findings from the continual learning literature which suggest catastrophic forgetting happens when two tasks are similar and therefore interfere (Farajtabar et al., 2020; Bennani et al., 2020; Doan et al., 2021). In general, non-factoid datasets see a lesser effect except GSM8K.